Add the two equations, 0.75x+1.5y=40 and -1.5x-1.5y=-60, to find the value of x.
Answer #1

soli. The given two equations 0.75*+1.54 - 40-0 -1.5% -1-5y =-60 - © Adding equo f eq 0+ 0.752 +1.59 = 40 -1.59 -1/5y = -60